I have answered this question but there are a few parts that I am struggling with

If the answer could be shown with workings it would be greatly appreciated

Complete the empty frames in the following table and show your work below (identify calculations by letter). For example, Spaling’s interest expense can be inferred from EBIT and Times interest earned, since TIE = EBIT / Interest expense.

EBIT = 300,000, TIE = 30, so 30 = 300,000 / Interest expense, and Interest expense = 300,000 / 30 = 10,000

Spaling

Preston

EBIT (Earnings before interest and taxes)

300,000

190,000

Interest expense

A

15,000

Net income

200,000

J

Dividend payout ratio

35%

K

Retention ratio

B

60%

Dividends declared during the year

C

40,000

Sales

3,000,000

L

Average assets during the year

D

1,500,000

Average debt during the year

700,000

M

Average shareholders’ equity during the year

1,800,000

N

Asset turnover ratio

E

1.3333

Debt ratio

F

0.3333=1/3

Return on sales

G

0.095

Return on assets

0.08

O

Return on equity

H

0.10

Market price per share, beginning of year

20

P

Market price per share, end of year

15

20

Total shareholder return

I

0.1556

Number of shares outstanding

150,000

50,000

Times interest earned

30

Q
Answer #1

A is already calculated.

B = retention ratio.

Retention ratio = 1 - dividend payout ratio = 1 - 35% = 1-0.35 = 0.65 or 65%

C = dividends declared during the year.

dividends declared during the year = dividend payout ratio * net income = 0.35* $200,000 = $70,000

D = Average assets during the year

Return on assets = Net income/ average assets during the year

Hence, average assets during the year = Net income/return on assets = $200,000/ 0.08 = $2,500,000

E = asset turnover ratio

= Revenue/total assets = $3,000,000/$2,500,000 = 1.2

F = debt ratio

= debt or liabilities/ total assets = $700,000/$2,500,000 = 0.28 or 28%

G = Return on sales

= operating profit/ net sales

operating profit = EBIT. Hence, return on sales = $300,000/$3,000,000 = 0.10 or 10%

H = return on equity = net income/ shareholders equity = net income / (assets-liabilities)

= $200,000 / ($2,500,000 - $700,000) = $200,000/$1,800,000 = 0.1111 = 11.11%

I = shareholders return

= [(market price of the share at the end of the year/ market price of the share at the beginning of the year)-1]*100

=($15/$20-1)*100 = -25%

K = dividend payout ratio = 1- retention ratio = 1-60% = 1-0.60 = 0.40 or 40%

J = net income. dividend payout ratio = dividends during the year/ net income

therefore, net income = dividends during the year/dividend payout ratio = $40,000/ 0.40 = $100,000

L = sales

Asset turnover ratio = Sales or revenue/total assets

Hence Sales = Asset turnover ratio*total assets = $1,500,000*1.3333 = $2,000,000

M = average debt during the year

debt ratio = average debt during the year/ total assets

hence, debt = debt ratio * total assets = 1/3 * $1,500,000 = $500,000

N = Average shareholders’ equity during the year

return on equity = net income/ shareholders equity

shareholders equity = net income/return on equity = $100,000/0.1 = $1,000,000

or shareholders' equity = (assets-liabilities) = $1,500,000- $500,000 = $1,000,000

O = return on assets

= Net income/ average assets during the year = $100,000/$1,500,000 = 0.06667 or 6.67%

P = market price at the beginning of the year

total shareholder return = (market price of the share at the end of the year/ market price of the share at the beginning of the year)-1

0.1556 = ($20 / market price of the share at the beginning of the year)-1

1.1556 * market price of the share at the beginning of the year = $20

Hence, market price of the share at the beginning of the year =$20/1.1556 = $17.31

Q = times interest earned = EBIT/ interest expense $190,000/$15,000 =12.667